OpenStudy (anonymous):

PLEASEEE HELP !! Evaluate the square root of one hundred plus two squared minus seven minus the square root of thirty-six. negative eight point nine 1 seventeen twenty-nine

OpenStudy (anonymous):

\[\sqrt{100}+2^{2}-7-\sqrt{36}=x\]

OpenStudy (anonymous):

You have to write out the expression, using your knowledge of mathematical sentences.

OpenStudy (anonymous):

Then solve accordingly. But for more in depth understanding consider this: square root of 100 means this:\[\sqrt{100}\] Does this help?